When a 500-mg medical dosage has degenerated to approximately 63 mg of its original, how many half-lives have passed?
Answer #1

Let n be number of half lives

Then:

N/No = (1/2)^n

63/500 = (1/2)^n

1/8 = (1/2)^n

(1/2)^3 = (1/2)^n

n = 3

Answer: 3